OpenStudy (anonymous):

can you please help me with this => Simplify Remember to use absolute-value notation when necessary. If a root cannot be simplified, state this. 1976 sqrt sign then (2a + b)^1976

OpenStudy (anonymous):

\[\sqrt[1976]{(2a+b)^{1976}}=|2a+b|\]

OpenStudy (anonymous):

how did you work this out please?

OpenStudy (anonymous):

what steps did you do to get the answer please help me, I am totally confused

OpenStudy (anonymous):

\[\sqrt[n]{x^n}=|x|\] if \(n\) is even

OpenStudy (anonymous):

for example \[\sqrt{(-3)^2}=\sqrt{9}=3\] or \[\sqrt[4]{(-2)^4}=\sqrt[4]{16}=2\] etc